OpenStudy (anonymous):

Solve the compound inequality. x ≤ 2 or x ≥ 5

OpenStudy (jdoe0001):

I don't see anything to solve there \(\bf {\color{red}{ x}} \le 2 \quad or\quad {\color{red}{ x}} \ge 5\implies 5\le {\color{red}{ x}} \le 2\)